# Do MLLMs Really Understand the Charts?

### v1

@@ Introduction
In Figure 3, powerful models like GPT-4o and Qwen2.5-VL-7B either misinterpret the query or resort to factual hallucination. In contrast, our ChartReasoner, sculpted by the RFT strategy, demonstrates a flawless, step-by-step reasoning process. This case study provides a compelling visual proof of our quantitative findings: ChartReasoner enhanced by the RFT strategy is transformed from a system prone to errors and hallucination into a reliable and structured visual reasoner that moves beyond merely optimizing for the answer.
Evaluation of ChartReasoner on the Public Benchmarks
The robust generalization of ChartReasoner is confirmed by its strong performance across a suite of public benchmarks, as detailed in Table 3. This success stems from our model’s core visual reasoning capability, honed through our training strategy, which is highly transferable to other reasoning-intensive tasks. This approach contrasts with that of other chart-specific models relying on SFT. Although SFT can improve performance on specific tasks, it often fails to instill a generalizable reasoning faculty, limiting extensibility. In contrast, ChartReasoner, enhanced by the RFT strategy, is specifically designed to cultivate this fundamental reasoning, validating the superiority of our method.
Conclusion
@@
In this paper, we investigate a compelling yet significant question: ”Do MLLMs really understand the charts?” By establishing the CRBench, we extensively evaluated open-source, close-source, and chart-specific MLLMs. The results shows a significant degradation in the performance of these models, and, through chain-of-thought reasoning, revealed their inability to estimate numerical values through visual reasoning, similar to human behavior. To address this issue, we propose ChartReasoner, which enhances the visual reasoning ability of MLLMs via an RFT strategy. This strategy first activates reasoning capabilities through SFT, and then generalizes reasoning abilities through RL. Extensive experiments conducted on the proposed CRBench and public chart reasoning datasets demonstrate the effectiveness of ChartReasoner. This work paves the way for empowering MLLMs to really understand the charts in a human-like manner.
@@
  • Sec. A provides details of the proposed benchmark, CRBench, including the task definition, the data construction pipeline, and data sources, topics, and chart types.
  • Sec. B describes the training strategy of the ChartReasoner, detailing the construction of the SFT and RL datasets.
  • Sec. C presents further details on the inference.
  • Sec. D includes other relevant details and extended results to supplement the main findings.
Appendix A A. CRBench Benchmark Details
Problem Definition
The primary task addressed by our benchmark, CRBench, is numerical value estimation on non-annotated charts. Formally, given a chart image $C$ and a query $Q$ that specifies a target data point within the chart, the goal is to produce a numerical answer $A$ that accurately estimates the value of that data point. Crucially, the chart image $C$ is non-annotated, meaning that the numerical values corresponding to graphical elements (e.g., the height of a bar, a point on a line) are not present as explicit text labels.
@@
Figure 4: The automated data generation pipeline for CRBench.
Data Generation Pipeline
The data generation pipeline for CRBench and training datasets, visualized in Figure 4, is an automated, multi-stage process designed to programmatically create a large and diverse set of non-annotated charts. The entire workflow is built upon generating and refining executable Python plotting scripts, ensuring that each data sample is verifiably correct.
@@
    Chart Generation by Large Language Model (LLM): The process begins with a curated seed pool of high-quality plotting scripts that serve as few-shot demonstrations. Using a Self-Instruct methodology, an LLM is then prompted to programmatically generate a vast and diverse set of new, non-annotated chart plotting scripts across 38 distinct topics.
@@
    Validation and Evolutionary Enhancement: Each generated script is rigorously validated via sandboxed execution to ensure it runs error-free. A key feature of the pipeline is an AI-driven repair loop, where the language model automatically debugs and corrects any code that fails execution. Subsequently, a portion of these validated scripts undergo an optional Evol-Instruct phase to further increase their visual and structural complexity.
@@
    Final Curation: Only scripts that successfully pass the entire validation and enhancement pipeline are accepted into the final datasets. These validated scripts are then executed to generate the final, high-fidelity chart images used in our benchmark and datasets.
Ground-Truth Q&A Generation and Verification.
After generating the chart images, we create the corresponding ground-truth Q&A pairs. We first use a LLM to generate questions and answers based on the chart’s source code. Each generated Q&A pair then undergoes a rigorous, multi-tiered verification: it is first cross-checked for accuracy by another LLM instance, and subsequently undergoes a final manual review by human annotators to guarantee its unequivocal correctness.
Data topics and Chart Examples
To ensure the breadth relevance of CRBench, the data generation process samples from a diverse set of 38 distinct topics, span a wide range of domains including finance, healthcare, technology, and social sciences, guaranteeing that the charts cover a variety of contexts and narratives. Furthermore, to robustly evaluate a model’s visual reasoning capabilities across different graphical representations, CRBench incorporates seven primary chart types. Figure 6 provides a representative example for each of these types, showcasing the visual diversity and complexity present in our benchmark.
Real-World Chart Collection
To complement our synthetic dataset and ensure CRBench reflects the challenges of real-world applications, we curate a substantial collection of charts from public online sources. This process involves careful sourcing from key data providers, a rigorous manual filtering stage, and a hybrid human-AI approach to generate high-quality question-answer pairs.
Statista
A significant portion of the real-world charts is sourced from Statista [^2] https://www.statista.com/, a leading global platform specializing in market and consumer data. Statista provides professional, data-driven visualizations for a wide array of industries, covering topics from economic indicators and market forecasts to technology trends and consumer behavior. The charts on this platform are typically clean, well-structured, and designed for clear communication in business and academic contexts.
The World in Data
The second major source is Our World in Data [^3] https://ourworldindata.org/, a renowned scientific online publication based at the University of Oxford. Its mission is to make data and research on the world’s most significant challenges, such as global health, economic development, and environmental change, accessible and understandable. The visualizations from the World in Data are often more complex and data-rich, including long-term time-series line charts, multi-country scatter plots, and stacked area charts.
In addition to these two primary repositories, the collection is supplemented by charts from various other miscellaneous public reports and online publications. All data collection is conducted in strict adherence to the copyright policies, terms of service, and licensing agreements of each source to ensure full ethical compliance.
Manual Filtering and Vetting.
A large pool of charts is initially collected. Each chart then undergoes a meticulous manual vetting process conducted by our team of annotators to ensure it met the stringent criteria for inclusion in CRBench. A chart is only accepted if it satisfied all of the following conditions:
@@
    High Visual Quality: The image has to be of sufficient resolution, clear, and free of significant compression artifacts or other visual noise that could impede interpretation.
@@
    Data Integrity: The chart has to be coherent and visually consistent, with clearly defined axes, legends, and graphical elements.
  3. 3.
    Absence of Annotations: We exclusively select charts where numerical values are not explicitly printed on the graphical elements (e.g., no numbers on top of bars or next to points on a line). This constraint is fundamental to our benchmark, as it forces a model to perform genuine visual reasoning rather than relying on OCR shortcuts.
Charts that fails to meet any of these standards are discarded from the final dataset.
Question-Answer Pair Generation.
Once a real-world chart is approved, we employ a two-stage, human-in-the-loop process to generate its corresponding question-answer pair:
@@
    MLLM-based Candidate Generation: We use a capable Multimodal Large Language Model (MLLM) to generate an initial set of candidate question-answer pairs for each chart. The model is prompted to ask a specific numerical estimation question based on the visual information in the chart.
  2. 2.
    Human Verification and Refinement: Every MLLM-generated pair is subsequently subjected to rigorous human review. Annotators would first verify the question’s clarity and relevance. Then, they would carefully perform the visual estimation themselves to validate the accuracy of the answer. If necessary, they would refine the question’s phrasing or correct the ground-truth answer to ensure the final Q&A pair was unambiguous and factually correct.
This careful, multi-stage curation process ensures that the real-world portion of CRBench is not only challenging and relevant but also contains reliable and accurate ground-truth data. Figure 7 showcases several examples of the final curated real-world charts from our collection, along with their corresponding question-answer pairs.
Appendix B B. Details of Datasets and Training
CoT Data Construction for Reasoning Activation
The dataset used for the Supervised Fine-Tuning (SFT) ”cold start” phase is meticulously constructed through a knowledge distillation process. The goal is to generate high-quality Chain-of-Thought (CoT) data that can effectively activate the reasoning paradigm of our base models. This process leverages a powerful teacher model(In our experiment, we use the Qwen2.5-VL-32B-Instruct as the teacher mdoel) to generate reasoning traces for the question-answer pairs in our CRBench dataset.
The data construction pipeline, implemented in the generation scripts, involves several key stages to ensure the quality and validity of the final CoT samples:
  1. 1.
    Prompting for Chain-of-Thought Generation: For each (C,Q,A) triplet from original data, we prompt the teacher model to generate a step-by-step reasoning process. Critically, the prompt(in Figure8) is designed to guide the model to derive the provided ground-truth answer, forcing it to articulate a logical pathway that leads to the correct result. The model is required to structure its output using specific tags, separating the reasoning steps (<think>...</think>) from the final answer (<answer>...</answer>).
@@
    • Answer Verification: The final answer extracted from the <answer> tag is programmatically compared against the ground-truth answer. We employ a robust evaluation function that checks for both exact string matches and numerical equivalence within a tolerance threshold to ensure correctness.
@@
This meticulous, multi-stage distillation and verification process—with examples showcased in Figure 9 and Figure 10—ensures that the data used for the reasoning activation phase is both correct and logically sound, providing a strong foundation for the subsequent reinforcement learning stage.
Dataset Construction for GRPO
The dataset for the Generative Rejection-based Policy Optimization (GRPO) stage is meticulously curated to focus the training on the most informative examples. Instead of using a random subsample of data, we construct a specialized dataset of moderately difficult samples. These are question-answer pairs that the base model can sometimes answer correctly but frequently gets wrong, representing the frontier of its reasoning capabilities and thus offering the richest signal for reinforcement learning.
The curation process inspired by the principles of rejection sampling and active learning, involves a multi-round, varied-prompting inference pipeline designed to probe the model’s knowledge boundaries. We first run inference on the entire CRBench dataset multiple times using the same base model. To elicit diverse responses, we systematically vary the prompts between rounds (e.g., alternating between direct answering and Chain-of-Thought prompts) and increase the model’s sampling temperature to 0.9. The correctness of every generated response is logged. After the evaluation is complete, we filter this log to isolate the key samples: we select only those question-answer pairs that the model answered correctly in at least one round but incorrectly in at least one other round.
@@
The rationale behind this selective filtering strategy is to maximize the efficiency and impact of the subsequent RL training. By focusing on these ”boundary” cases, we avoid wasting computational resources on problems the model has already mastered (always correct) or those that are currently far beyond its reach (always incorrect). Training on these ”boundary” samples forces the policy to learn to distinguish between successful and flawed reasoning pathways for the exact same problem. This targeted approach ensures that the GRPO stage is dedicated to resolving ambiguity and reinforcing robust reasoning where it is most needed, leading to more significant and generalizable improvements in the model’s core reasoning abilities.
Detailed Formulation of the Accuracy Reward
@@
A core component of the GRPO framework is the Accuracy Reward ($R_{\text{acc}}$), which is designed to provide a dense, fine-grained signal for optimizing the model’s numerical estimation capabilities. A simple binary reward (correct/incorrect) is often too sparse for reinforcement learning, as it fails to differentiate between a near-miss and a completely wrong answer. To overcome this, we designed a continuous reward function that recognizes and rewards ”nearly correct” answers, thereby creating a smoother optimization landscape. The formulation of this reward is visualized in Figure 5.
Figure 5: Visualization of our quadratic accuracy reward function ($R_{\text{acc}}$) with a tolerance threshold of $\varepsilon=0.02$. The reward smoothly decays from 1 to 0, with a gentle slope for small errors and an accelerating penalty for larger deviations.
For a predicted numerical answer $A_{\text{pred}}$ and a non-zero ground-truth answer $A_{\text{gt}}$, we first calculate the relative error:
Based on a tolerance threshold of $\varepsilon=0.02$, we define the reward using a piecewise quadratic function that smoothly decays from a maximum of 1 to 0:
$R_{\text{acc}}(d_{\text{rel}})=\begin{cases}\left(1-\frac{d_{\text{rel}}}{\varepsilon}\right)^{2}&\text{if }d_{\text{rel}}<\varepsilon\\ 0&\text{if }d_{\text{rel}}\geq\varepsilon\end{cases}$ (7)
@@
Appendix C C. Inference Details
This section provides additional details regarding our inference and evaluation setup to ensure full reproducibility.
Inference Hyperparameters and Framework
All experiments reported in this paper were conducted using the default hyperparameters of each respective model, with no model-specific tuning performed at inference time. This ensures a fair and consistent comparison across all models. To guarantee reproducibility, the random seed for all generation processes was fixed to 42. Our inference pipeline is built upon the vLLM framework, which provides efficient, high-throughput serving for Large Language Models.
Prompting Strategies for Evaluation
To thoroughly investigate the models’ reasoning capabilities and behaviors, we employed three distinct prompting strategies during evaluation, designed to control the externalization of the model’s thought process. The first is Direct Answering, a baseline prompt that instructs the model to provide only the final answer. The second, Optional Chain-of-Thought (Optional-CoT), encourages but does not require the model to generate a reasoning trace, allowing us to observe its natural tendency to reason. The third, Forced Chain-of-Thought (Forced-CoT), strictly requires a step-by-step reasoning process before the final answer, compelling the model to articulate its logic. As illustrated in Figure 14 and Figure 15, which showcases two comparative examples, these varied strategies allow us to gain deeper insights into not just what the models answer, but how their performance is affected by the explicit demand for reasoning.
Appendix D D. Evaluation Protocols for Specialized Models and Benchmarks
This section outlines the precise met hodologies used to evaluate specialized, chart-specific models and to benchmark our ChartReasoner on existing public datasets. Our protocols were designed to ensure maximum fairness and comparability by strictly adhering to the author-recommended settings and official evaluation scripts for each model and benchmark.
Evaluation of Chart-Specific Models on CRBench
Evaluating the performance of existing chart-specific models on our proposed CRBench required a careful approach to ensure a fair comparison. These models, including ChartGemma, TinyChart, ChartInstruct, ChartVLM, and ChartLlama, are often fine-tuned on specific data formats and are highly optimized for a particular style of prompting.
For this reason, we did not use our generalized prompting templates (as shown in Figures 12 and 13) for these models. Instead, to elicit the best possible performance from each model and establish the strongest baseline, we followed this procedure:

### v2

@@ 5.4 Case Study
In Figure 3, powerful models like GPT-4o and Qwen2.5-VL-7B either misinterpret the query or resort to factual hallucination. In contrast, our ChartVR, sculpted by the RFT strategy, demonstrates a flawless, step-by-step reasoning process. This case study provides a compelling visual proof of our quantitative findings: ChartVR enhanced by the RFT strategy is transformed from a system prone to errors and hallucination into a reliable and structured visual reasoner that moves beyond merely optimizing for the answer.
6 Conclusion
@@
In this paper, we investigate a compelling yet significant question: ”Do MLLMs really understand the charts?” By establishing the ChartVRBench, we extensively evaluated open-source, close-source, and chart-specific MLLMs. The results shows a significant degradation in the performance of these models, and, through chain-of-thought reasoning, revealed their inability to estimate numerical values through visual reasoning, similar to human behavior. To address this issue, we propose ChartVR, which enhances the visual reasoning ability of MLLMs via an RFT strategy. This strategy first activates reasoning capabilities through SFT, and then generalizes reasoning abilities through RL. Extensive experiments conducted on the proposed ChartVRBench and public chart reasoning datasets demonstrate the effectiveness of ChartVR. This work paves the way for empowering MLLMs to really understand the charts in a human-like manner.
@@
  • Sec. A provides details of the proposed benchmark, ChartVRBench, including the problem definition, data sources, topics, and chart types.
  • Sec. B describes the training strategy of the ChartVR, detailing the construction of the SFT and RL datasets.
  • Sec. C presents further details on the evaluation and inference.
  • Sec. D interpret the human performence in ChartVRBench.
Appendix A ChartVRBench Benchmark Details
A.1 Problem Definition
The primary task addressed by our benchmark, ChartVRBench, is numerical value estimation on non-annotated charts. Formally, given a chart image $C$ and a query $Q$ that specifies a target data point within the chart, the goal is to produce a numerical answer $A$ that accurately estimates the value of that data point. Crucially, the chart image $C$ is non-annotated, meaning that the numerical values corresponding to graphical elements (e.g., the height of a bar, a point on a line) are not present as explicit text labels.
@@ A.1 Problem Definition
A.2 Data topics and Chart Examples
Table 6: The 38 topics covered in the ChartVRBench dataset.
Category Category
Agriculture and Food Production Human Resources and Employee Management
Architecture and Building Language and Communication
Artificial Intelligence and Robotics Law and Legal Affairs
Art and Design Literature and Writing
Astronomy and Space Manufacturing and Production
Biology and Life Sciences Marketing and Advertising
Books and Publishing Mathematics and Statistics
Business and Finance Music and Performance
Computer Science and Information Technology Physics and Chemistry
Education and Academics Real Estate and Housing Market
Energy and Utilities Religion and Spirituality
Environment and Sustainability Retail and E-commerce
Fashion and Style Science and Engineering
Film and Cinema Social Media and the Web
Food and Beverage Industry Social Sciences and Humanities
Futurism and Innovation Society and Community
Government and Public Policy Sports and Entertainment
Healthcare and Health Transportation and Logistics
History and Culture Travel and Exploration
To ensure the breadth relevance of ChartVRBench, the synthetic data generation process samples from a diverse set of 38 distinct topics, as shown in Table 6 guaranteeing that the charts cover a variety of contexts and narratives. Furthermore, to robustly evaluate a model’s visual reasoning capabilities across different graphical representations, ChartVRBench incorporates seven primary chart types. Figure 5 provides a representative example for each of these types, showcasing the visual diversity and complexity present in our benchmark.
A.3 Real-World Chart Collection
To ensure ChartVRBench reflects the challenges of real-world applications, we curated a substantial collection of charts from public online sources. This section details our three-stage process: data sourcing, a rigorous manual filtering protocol, and a hybrid human-AI pipeline for generating high-quality question-answer pairs. Figure 6 showcases several examples of the final curated real-world charts from our collection.
A.3.1 Data Sourcing
Statista.
A significant portion of the real-world charts was sourced from Statista [^2] https://www.statista.com/, a leading global platform specializing in market and consumer data. Statista provides professional, data-driven visualizations for a wide array of industries, covering topics from economic indicators and market forecasts to technology trends and consumer behavior.
Our World in Data.
The second major source was Our World in Data [^3] https://ourworldindata.org/, a renowned scientific online publication based at the University of Oxford. Its mission is to make data and research on the world’s most significant challenges, such as global health, economic development, and environmental change, accessible and understandable through complex and data-rich visualizations.
In addition to these two primary repositories, the collection was supplemented by charts from various other miscellaneous public reports and online publications. All data collection was conducted in strict adherence to the copyright policies, terms of service, and licensing agreements of each source to ensure full ethical compliance.
A.3.2 Curation Protocol
Once a large pool of charts was collected, each candidate chart underwent a meticulous, two-step curation process performed by our recruited team of human annotators.
Step 1: Manual Filtering and Vetting.
Each chart was manually vetted against stringent criteria for inclusion in ChartVRBench. A chart was accepted only if it satisfied all of the following conditions, otherwise it was discarded:
@@ A.3.2 Curation Protocol
    High Visual Quality: The image must be of sufficient resolution, clear, and free of significant compression artifacts or other visual noise that could impede interpretation.
@@
    Data Integrity: The chart must be coherent and visually consistent, with clearly defined axes, legends, and graphical elements.
@@
    Absence of Annotations: We exclusively select charts where numerical values are not explicitly printed on the graphical elements (e.g., no numbers on top of bars). This constraint is fundamental to our benchmark, as it forces a model to perform genuine visual reasoning rather than relying on OCR shortcuts.
Step 2: Question-Answer Pair Generation.
Once a chart was approved, we employed a two-stage, human-in-the-loop process to generate its corresponding question-answer pair:
@@
    MLLM-based Candidate Generation: We first use a capable MLLM to generate an initial set of candidate question-answer pairs for each chart, prompting it to ask a specific numerical estimation question.
@@
    Human Verification and Refinement: Every MLLM-generated pair is then subjected to rigorous human review. Annotators verify the question’s clarity and relevance, and then carefully perform the visual estimation themselves to validate the answer’s accuracy. If necessary, they refine the question’s phrasing or correct the ground-truth answer to ensure the final Q&A pair is unambiguous and factually correct.
Appendix B Details of Datasets and Training
B.1 Data Construction for SFT
The dataset used for the Supervised Fine-Tuning (SFT) “cold start” phase is meticulously constructed through a knowledge distillation process. The goal is to generate high-quality Chain-of-Thought (CoT) data that can effectively activate the reasoning paradigm of our base models.
While we utilize the same underlying generation pipeline (rendering engine and topic distribution) to ensure domain consistency, the specific chart instances and question-answer pairs in the SFT set are distinct from those in the benchmark. Crucially, to maintain strict train-test separation, this SFT dataset is generated as a completely independent batch from the ChartVRBench evaluation set.
This process leverages a powerful teacher model (Qwen2.5-VL-32B-Instruct) to generate reasoning traces for this training-specific corpus. The data construction pipeline involves several key stages to ensure the quality and validity of the final CoT samples:
@@ B.1 Data Construction for SFT
    Prompting for Chain-of-Thought Generation: For each generated training instance, consisting of Python plotting code (C) and a question (Q), we employ the teacher LLM to generate a step-by-step reasoning process. The core of this process involves a carefully constructed textual prompt that integrates the Python code (C) and the question (Q). This prompt is designed to force the model to articulate a logical pathway from the code and question to the correct result. The model is required to structure its output using specific tags, separating the reasoning steps (<think>...</think>) from the final answer (<answer>...</answer>).
@@
    • Answer Verification: The final answer extracted from the <answer> tag is programmatically compared against the ground-truth answer derived from the code. We employ a robust evaluation function that checks for both exact string matches and numerical equivalence within a tolerance threshold to ensure correctness.
@@
B.2 RL Algorithm Selection
We selected GRPO to fine-tune our multimodal model for enhanced chart visual reasoning, primarily due to its superior efficiency and its alignment with our reward structure. Compared to Proximal Policy Optimization (PPO), GRPO significantly reduces computational and memory overhead. GRPO eliminates the need for a separate value model—which is typically as large as the policy model—by estimating the baseline directly from the scores of multiple sampled outputs. This efficiency is critical given the large scale of our models (e.g., 7B parameters), making GRPO a practical solution under limited hardware resources.
Furthermore, while Direct Preference Optimization (DPO) offers an efficient alternative to traditional RLHF, it is fundamentally designed for binary preference datasets (i.e., chosen vs. rejected responses). Our task, however, benefits from a more granular, continuous reward signal that reflects the degree of correctness in quantitative analysis. GRPO is adept at directly optimizing for such programmatic, scalar rewards, allowing the model to learn from fine-grained feedback. This makes it better suited for improving the visual reasoning in chart than a preference-based method like DPO.
B.3 Dataset Construction for GRPO
The curation process, inspired by the principles of rejection sampling and active learning, involves a multi-round, varied-prompting inference pipeline designed to probe the model’s knowledge boundaries. The goal of this pipeline is to construct a specialized, high-signal dataset by isolating ambiguous cases that the SFT-tuned model can sometimes solve but not consistently. This strategy focuses the training process on the most informative examples where the model is most uncertain, rather than wasting computational resources on problems that are already mastered (always correct) or are currently too difficult (always incorrect).
Our pipeline involves the following systematic steps:
  1. 1.
    Initial Dataset Curation: We begin by constructing an initial, high-quality dataset for GRPO. This dataset is synthesized following the method introduced earlier. We selected only those instances that achieved a score of either 4 or 5, ensuring a strong baseline of correct.
  2. 2.
    Multi-Round Inference: Initially, we run inference multiple times on the initial GRPO dataset using our base model. To elicit a wide range of reasoning pathways and outcomes for each problem, we set the sampling temperature to 1.0 for each run.
  3. 3.
    Filtering for ”Stochastic Correctness”: The correctness of every generated response is logged. After all rounds are complete, we filter this log to isolate the target samples. We select only those question-answer pairs that the model answered correctly in at least one round but incorrectly in at least one other round.
@@ B.3 Dataset Construction for GRPO
The rationale behind this selective filtering is to force the policy to learn to distinguish between successful and flawed reasoning pathways for the exact same problem. Training on these ”boundary” cases ensures that the GRPO stage is dedicated to resolving ambiguity and reinforcing robust reasoning where it is most needed, leading to more significant and generalizable improvements in the model’s core abilities.
B.4 Detailed Formulation of the Accuracy Reward
@@
A core component of the GRPO framework is the Continuous Accuracy Reward ($R_{\text{acc}}$), which is designed to provide a dense, fine-grained signal for optimizing the model’s numerical estimation capabilities. A simple binary reward (correct/incorrect) is often too sparse for reinforcement learning, as it fails to differentiate between a near-miss and a completely wrong answer. To overcome this, we designed a continuous reward function that recognizes and rewards ”nearly correct” answers, thereby creating a smoother optimization landscape.
@@
B.5 More Training Details
Our training process begins with the Qwen2.5-VL-7B-Instruct model as the foundation. Using the Swift framework, we perform SFT for 2 epochs on our 4.2k instruction-following dataset. In this stage we freeze the vision tower and the multimodal aligner while exclusively tuning the LLM’s parameters. We set the learning rate to 1e-5 with a warm-up ratio of 0.05 and use an effective batch size of 256. The SFT process is conducted on 8 NVIDIA H800 GPUs, utilizing bfloat16 precision and the DeepSpeed ZeRO-3 optimization strategy.
For the GRPO stage, we initialize the model with the checkpoint from the SFT phase and employ the GRPO algorithm on our 3.4k preference dataset. In this phase, we continue to freeze the vision tower but expand the scope of fine-tuning to include both the LLM and the multimodal aligner. The learning rate is reduced to 1e-6, again with a 0.05 warm-up ratio. For the rollout process, we use a generation batch size of 32 to create 4 completions per sample with a temperature of 1.0; the training itself uses an effective batch size of 64. The reward function is the composite of the Format and Continuous Accuracy rewards. The hardware and optimization setup remains consistent, utilizing 8 NVIDIA H800 GPUs with bfloat16 precision and DeepSpeed ZeRO-3.
Appendix C Evaluation and Inference Details
This section outlines the precise methodologies and inference settings used to evaluate all models and benchmarks, ensuring full reproducibility and fairness. Our protocols were designed by strictly adhering to the author-recommended settings and official evaluation scripts where available.
C.1 General Inference Settings
All experiments reported in this paper were conducted using the default hyperparameters of each respective model, with no model-specific tuning performed at inference time. To ensure reproducibility, the random seed for all generation processes was fixed to 42, and the sampling temperature was set to 1.0. Our inference pipeline is built upon the vLLM framework, which provides efficient, high-throughput serving for Large Language Models.
C.2 Evaluation on ChartVRBench
Our evaluation on the proposed ChartVRBench employed different prompting strategies depending on the model type to ensure a fair and rigorous assessment.
General MLLMs.
To verify the visual reasoning capabilities of general-purpose models, we employed a structured Chain-of-Thought (CoT) prompt, shown in Figure 10. This prompt compels the model to first articulate its reasoning—by analyzing axes, data points, and context—before providing a final answer. The prompt enforces a strict separation between the step-by-step logic (output in ‘¡think¿‘ tags) and the concise final output (in an ‘¡answer¿‘ tag). This approach allows us to pinpoint the exact stage where a model’s logic succeeds or fails, moving our analysis beyond simple accuracy metrics.
Chart-Specific Models.
In contrast, for models already fine-tuned on specific chart-related data formats (including ChartGemma, TinyChart, ChartInstruct, ChartVLM, and ChartLlama), we did not use our generalized CoT prompt. To elicit their best possible performance and establish the strongest baseline, we followed the official author-recommended procedures: